# Reinforcement Unlearning

### v2

@@ 3.4. Comparison of the Two Methods
One advantage of the poisoning-based method over the decremental reinforcement learning-based method is its ability to address the over-unlearning issue associated with the latter. The decremental reinforcement learning-based method may inadvertently suffer from over-unlearning, which occurs when the deep reinforcement learning model is fine-tuned to degrade the agent’s learning performance in $\mathcal{M}_{u}$. This issue is also observed in conventional machine unlearning scenarios (Bourtoule et al., 2021). Even if efforts are made to restrict the deterioration to $\mathcal{M}_{u}$, it may still affect other environments due to the shared distribution among them. However, the poisoning-based method inherently avoids this issue by focusing on enabling the agent to learn new knowledge rather than intentionally forgetting existing knowledge. Thus, the poisoning-based method has the potential to achieve superior performance in non-unlearning environments compared to the decremental RL-based method.
@@
A valid concern regarding the poisoning-based method is the potential occurrence of catastrophic forgetting, which arises when the continual updating of the deep reinforcement learning model results in the overwriting of previously acquired knowledge. However, this issue does not arise in the context of the poisoning-based method. The primary cause of catastrophic forgetting is a shift in the input distribution across different environments (Kirkpatrick and et al., 2017; Shi et al., 2021). In our scenario, the modified environment $\mathcal{M}^{\prime}_{u}$ retains the same distribution as the other environments. This is because the modification is solely applied to the transition function, while the state and action spaces, as well as the reward function, remain unchanged. Specifically, the transition function dictates the evolution of states based on the actions taken by the agent, governing how states change over time. In contrast, the state and action spaces, as well as the reward function, are pre-defined by the model owner during the training of the agent. These foundational elements remain relatively unaffected by modifications to the transition function. Thus, there is no distribution shift across environments in our problem, thereby mitigating the risk of catastrophic forgetting.
3.5. Environment Inference
@@
One of our contributions is to propose a new evaluation methodology named environment inference. This kind of inference aims to infer an agent’s training environments by observing agent’s behavior (Pan et al., 2019). By employing this approach, we can assess whether the agent has successfully eradicated the knowledge of the unlearning environment. If the removal of the unlearning environment’s knowledge is executed correctly, the agent’s behavior in that environment should be random rather than purposeful. Thus, by observing the agent’s behavior, an adversary can only infer a randomized environment, devoid of any specific knowledge.
A notable environment inference (Pan et al., 2019) utilizes a genetic algorithm to identify a transition function that not only satisfies specific constraints but also provides the best possible explanation for the observed policy. Inspired by this approach, our research also employs the same genetic algorithm to infer the unlearning environments. Specifically, by observing the agent’s behavior, we can obtain a policy $\pi_{target}$. In each iteration, we maintain a set of dynamic transition functions. For each transition function $\mathcal{T}$, we train an optimal policy $\pi^{*}_{\mathcal{T}}$. By quantifying the similarity between $\pi_{target}$ and $\pi^{*}_{\mathcal{T}}$, we obtain a fitness score denoted as $Score(\pi_{target},\pi^{*}_{\mathcal{T}})$. Our objective is to identify a transition function $\mathcal{T}$ that maximizes this fitness score. To achieve this objective, the top $n$ candidates are selected based on their fitness scores and carried over to the next generation. The remaining candidates are generated by involving two parents selected from the previous generation, with the selection based on their scores. These selected parents undergo a two-point crossover, generating child candidates. Finally, to introduce diversity, a random mutation is applied to the child candidates.
4. Experimental Evaluation
4.1. Experimental Setup
@@
Evaluation metrics in conventional machine unlearning (Xu et al., 2023) are not applicable to reinforcement unlearning. For instance, in reinforcement unlearning, there are no specific datasets to be forgotten, rendering metrics like “accuracy on forget set” irrelevant. Thus, it is necessary to propose new metrics.
@@ 4.1.1. Tasks
The experiments were conducted across four learning tasks: grid world, aircraft landing, virtual home and maze explorer. The virtual home and maze explorer tasks were sourced from the VirtualHome (VirtualHome, 2023) and MazeExplorer (Harries et al., 2019), respectively, while the other two tasks were developed by us. Although there are well-known RL tasks available, e.g., Gym (Gym, 2023) and Atari (Artari, 2023), they were deemed unsuitable for our research as those RL tasks are designed for single-environment and do not support multiple environments.
@@
Grid World.  The objective for the agent in this task is to navigate towards the predetermined destination. This task mirrors numerous challenges present in real-world autonomous driving and navigation scenarios, with each square akin to a section of road, intersection, or obstacle. These obstacles can represent restricted areas in reality. Thus, to forget the restricted areas, the agent must be able to unlearn these obstacles within the environment.
Aircraft Landing.  This task simulates an aircraft landing on the ground by avoiding the obstacles. This task pertains to autonomous systems, operating in critical environments and avoiding dangers.
@@
Virtual Home.  Virtual home is a multi-agent platform designed to simulate various activities within a household setting. This task mirrors real-world smart home challenges, with each room representing a different area of the home. Homeowners can request the agent to forget certain features, necessitating our reinforcement unlearning schemes for privacy maintenance.
@@
Maze Explorer.  Maze explorer is a customizable 3D platform. The objective is to guide an agent through a procedurally generated maze to collect a predetermined number of keys. This task represents scenarios where agents need to learn from visual information, making it relevant to applications such as robotic exploration in intricate environments.
@@
In these tasks, four available actions include moving up (forward), down (backward), left, and right. Environments are instantiated with predetermined sizes. The instantiation process involves two steps. Firstly, each environment is randomly generated, introducing variability in the placement of obstacles. Subsequently, a manual inspection is carried out to eliminate any instances of “dead locations”. These are locations within the environment that become inaccessible due to being entirely surrounded by obstacles.
@@
While our unlearning methods are evaluated in tasks with discrete state and action spaces, they can also be applied to tasks with continuous state and action spaces. Our approaches are independent of the underlying RL algorithms. To address continuous spaces, one can simply integrate our unlearning techniques with suitable RL algorithms, e.g., (Lee et al., 2018), designed for such environments.
4.1.2. Comparison Methods
@@
Learning From Scratch (LFS).  This method entails removing the unlearning environment and subsequently retraining the agent from scratch using the remaining environments when an unlearning request is received (Bourtoule et al., 2021). However, this method is not a desirable criterion for reinforcement unlearning. In the forthcoming experimental results, we will show that this approach fails to fulfill the objectives of reinforcement unlearning as defined in Section 3.1.
Non-transferable Learning From Scratch (Non-transfer LFS). To align with the objectives of reinforcement unlearning, we introduce a non-transferable learning-from-scratch approach. This approach is similar to the previously mentioned learning-from-scratch approach. However, a crucial distinction lies in the non-transferable version, which incorporates the non-transferable learning technique (Wang et al., 2022b) to restrict the approach’s generalization ability within the unlearning environments. In this approach, while training an agent, the model owner meticulously stores experience samples acquired from all learning environments, labeling them according to their source environment. When an unlearning request is initiated, the model owner engages in an offline retraining process. Specifically, all the collected experience samples are utilized in retraining the agent. If a sample originates from the unlearning environment, an inverse loss function is applied to minimize the agent’s cumulative reward. Conversely, for samples from other environments, the standard loss function is used to maximize the agent’s overall reward. Denoting the unlearning environment as $\mathcal{M}_{u}=(\mathcal{S}_{u},\mathcal{A}_{u},\mathcal{T}_{u},r)$, the loss functions are defined in Eq. 10.
(10) =$\mathcal{L}=\begin{cases}-\frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max\limits_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)\\ -Q(s_{t},a_{t};\theta))^{2}],&\text{if $s_{t}\in\mathcal{S}_{u}$},\\ \frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max\limits_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)\\ -Q(s_{t},a_{t};\theta))^{2}],&\text{otherwise}.\end{cases}$
Indeed, we have also conducted evaluations on a random-walking agent, which uniformly selects actions in each state. This random-walking agent functions as a representation of a completely unlearned agent, signifying it has forgotten all prior knowledge and returned to the initial status. This can serve as a foundational benchmark in the evaluation. However, the performance of this agent is notably degraded, requiring, for instance, several thousand steps to accomplish a given task. While these results provide valuable insights, their inclusion in the figures can obscure the clarity of other methods under consideration. Thus, to maintain the focus of the paper, we do not include these results in the final presentation.
4.1.3. Sample Complexity of Unlearning Methods
@@
There might be a concern regarding our proposed methods, as they allow the agent to engage in additional interactions with the unlearning environment. In contrast, both LFS and Non-transfer LFS do not involve further interactions with any environments. However, this additional interaction in our methods does not bring any extra advantages. The purpose of engaging with the unlearning environment is solely to collect experience examples. These experience samples are not required by either LFS or Non-transfer LFS since their objective is precisely to forget this information. Therefore, the absence of these samples, i.e., the lack of such interactions, does not impact the performance of both LFS and Non-transfer LFS.
4.2. Overall Performance
Before showing the final performance of the proposed unlearning methods, we first present the progression of their performance change during unlearning in grid world and aircraft landing.
@@ 4.2. Overall Performance
To illustrate the effectiveness of unlearning, we introduce the concept of “forget quality” as a quantitative measure of the strength of forgetting, aligning with the latest machine unlearning evaluation criteria (Maini et al., 2024) and tailoring it to reinforcement unlearning. We define the “cumulative truth ratio” to gauge this strength, serving as the foundation for the computation of forget quality. Specifically, the cumulative truth ratio $R_{truth}$ can be written as follows.
(11) $R_{truth}=\frac{\sum_{s\in\tau}\frac{1}{|\mathcal{A}_{w}|}\sum_{\hat{a}\in\mathcal{A}_{w}}\pi(\hat{a}|s)}{\sum_{s\in\tau}\pi(\tilde{a}|s)},$
@@
where $\hat{a}$ represents a wrong action from the wrong action set $\mathcal{A}_{w}$, $\tilde{a}$ is the correct action, and $\tau=((s_{1},a_{1}),\ldots,(s_{N},a_{N}))$ signifies a trajectory in $\mathcal{M}_{u}$. Here, the correct action is defined as the one capable of moving the agent closer to the target compared to other available actions in a given state, while the remaining actions are considered wrong. In particular, the cumulative truth ratio $R_{truth}$ quantifies the ratio between the average probability of selecting wrong actions and the probability of taking the correct action. Thus, a to-be-unlearned agent is expected to achieve a low $R_{truth}$, while an effectively unlearned agent in the unlearning environment should exhibit a high $R_{truth}$, resembling an agent that has never seen $\mathcal{M}_{u}$. For each agent, we obtain $N$ values of $R_{truth}$ each from the first $i$ ($i\in[1,N]$) step(s), and normalize them to serve as an empirical cumulative distribution function (ECDF). A two-sample Kolmogorov-Smirnov (KS) test (Massey Jr, 1951) is conducted between the ECDFs of the unlearned and retained agent to generate a $p$-value to quantify forget quality. A high $p$-value indicates a strong forgetting, implying that the $R_{truth}$ distributions of the unlearned agent and the retained agent are identical.
@@
Figure 4. Forget Quality vs. Model Utility. Larger marker represents more unlearning epochs.
@@
We compute the utility of the unlearned/retained model as the average cumulative rewards of it in the retaining environments. Figure 4 illustrates the trade-off between model utility and forget quality during unlearning in grid world and aircraft landing, where a larger marker denotes more unlearning epochs. In grid world (the left subfigure), the decremental RL-based method exhibits consistently high forget quality, while the poisoning-based method displays some fluctuations in forget quality but eventually attains a commendable level. This variance may be attributed to the randomness introduced by the RL algorithm in the poisoning-based approach. During the learning of the poisoning strategy, the method gradually converges, ultimately achieving a favorable result. Furthermore, both methods attain a higher forget quality than before unlearning, indicating the success of the unlearning process.
@@
For model utility, both methods demonstrate robust performance, fluctuating in a narrow range of $(25.075,25.175)$. This performance is comparable to that of the retained models trained by the two baseline methods, i.e., learning-from-scratch (LFS) and Non-transfer LFS. A closer examination of model utility reveals a dynamic shift in utility in both methods as the unleanring progresses. This oscillation may arise from two factors. First, the decremental RL-based method aims to erase the agent’s knowledge, potentially leading to over-forgetting (Hu et al., 2024). Second, the poisoning-based method involves an intricate interplay between the RL dynamics and the strategic introduction of poison. This interplay can introduce randomness.
A similar trend is observed in aircraft landing (the right subfigure). The primary distinction is that the poisoning-based method exhibits fewer fluctuations in forget quality in aircraft landing compared to grid world. This discrepancy may be attributed to the less complex nature of the aircraft landing setting, allowing the poisoning strategy to converge more rapidly.
Refer to caption (a) Average steps of the four methods Refer to caption (b) Rewards of the four methods
Figure 5. Four methods in the unlearning environment
Next, we make detailed comparisons between our methods and the baselines. In Figure 5, it becomes evident that the unlearning results of the LFS baseline method in all four experimental settings are subpar. The agent’s performance in the unlearning environment remains nearly unchanged before and after unlearning. The reason for this result lies in the agent’s ability to generalize knowledge from other environments and apply it to the unlearning environment, despite never having encountered it before.
@@
During training, the agent learns underlying rules and strategies from various environments. For instance, in the grid world setting, the agent acquires knowledge that obstacles should be avoided while collecting the target as quickly as possible. This learned knowledge, even if it was acquired in different environments, enables the agent to still perform well in unseen environments, including the unlearning environment. As a result, the baseline method proves to be ineffective as a reinforcement unlearning technique.
@@
4.3. Hyperparameter Study

### v5

@@ III-D Discussion of the Two Methods
Over-unlearning. The decremental RL-based method may inadvertently suffer from over-unlearning, which occurs when the model is fine-tuned to degrade the agent’s learning performance in $\mathcal{M}_{u}$. Even if efforts are made to restrict the deterioration to $\mathcal{M}_{u}$, it may still affect other environments due to the shared distribution among them. However, the poisoning-based method inherently avoids this issue by focusing on enabling the agent to learn new knowledge rather than intentionally forgetting existing knowledge. Thus, the poisoning-based method has the potential to achieve superior performance in non-unlearning environments compared to the decremental RL-based method.
Another factor contributing to the over-unlearning issue in the decremental RL-based method is the delayed influence of the second term in Eq. 5. This term is responsible for fine-tuning the policy to sustain performance in the retained environments. However, its impact may not manifest immediately. This delayed interference implies that the policy is tuned to initially focus solely on minimizing performance in the unlearning environment, without considering the impact on performance in the retained environments. As a result, the policy may prioritize forgetting the unlearning environment over preserving knowledge from the retained environments.
@@
Catastrophic Forgetting. A valid concern regarding the poisoning-based method is the potential occurrence of catastrophic forgetting, which arises when the continual updating of the model results in the overwriting of previously acquired knowledge. However, this issue does not arise in the context of the poisoning-based method. The primary cause of catastrophic forgetting is a shift in the input distribution across different environments [19, 33]. In our scenario, the modified environment $\mathcal{M}^{\prime}_{u}$ retains the same distribution as the other environments. This is because the modification is solely applied to the transition function, while the state and action spaces, as well as the reward function, remain unchanged. Specifically, the transition function dictates the evolution of states based on the actions taken by the agent, governing how states change over time. In contrast, the state and action spaces, as well as the reward function, are pre-defined by the model owner during the training of the agent. These foundational elements remain relatively unaffected by modifications to the transition function. Thus, there is no distribution shift across environments in our problem, thereby mitigating the risk of catastrophic forgetting.
Offline RL Settings. Although our methods are designed to online RL settings, they can be adapted to offline settings, where the agent cannot interact with the environment in real-time but relies on previously collected experience samples. The decremental RL-based method can utilize these samples by adjusting the policy to minimize performance in the unlearning environment while maintaining performance in retained environments. The poisoning-based method can modify the collected experience samples, e.g., states and rewards, to simulate the effect of altering the transition function.
Direct Reward Inversion. As both unlearning methods focus on reducing the agent’s received reward in the unlearning environment, a seemingly straightforward approach involves directly inverting the received reward: changing a real reward $r$ to $-r$. However, this method risks deteriorating the agent’s performance not only in the unlearning environment but also in remaining ones. This occurs because such inversion disrupts the entire reward structure foundational to the agent’s learning and decision-making processes, potentially leading to unintended behavior across all environments.
Mitigating Security Issues. The potential security issues associated with our unlearning methods can be effectively mitigated, particularly concerning adversarial attempts to reconstruct the unlearning environment by observing the agent’s behavior and leveraging knowledge that this behavior results from minimizing rewards. This is because both methods employ randomness in their action selection processes that goes beyond merely minimizing rewards. In the decremental RL-based method, while rewards are minimized, the policy also includes a degree of randomness in action selection that are not directly tied to reward signals. This randomness ensures that the behavior observed by an adversary does not consistently align with the most minimized rewards, thereby obscuring the true characteristics of the unlearning environment. Likewise, the poisoning-based method not only reduces reward, but also alters the transition dynamics between states. This means that even if an adversary can observe that certain actions are less rewarded, the underlying state transitions are altered. This alteration further obscures the patterns in the unlearning environment, complicating any adversary attempts to reconstruct the environment based on minimized rewards alone.
III-E Environment Inference
@@
One of our contributions is to propose a new evaluation methodology named environment inference. This kind of inference aims to infer an agent’s training environments by observing agent’s behavior [29]. By using this approach, we can assess if the agent has successfully erased the knowledge of the unlearning environment. If the removal of the unlearning environment’s knowledge is executed correctly, the agent’s behavior in that environment should be random rather than purposeful. Thus, by observing the agent’s behavior, an adversary can only infer a randomized environment.
A notable environment inference [29] utilizes a genetic algorithm to identify a transition function that not only satisfies specific constraints but also provides the best possible explanation for the observed policy. Inspired by this approach, our research also employs the same genetic algorithm to infer the unlearning environments. Initially, we randomly generate $L$ transition functions, representing the population. These transition functions undergo iterative updates through crossover and mutation operations. During the crossover process, pairs of transition functions are selected based on their fitness, determined by their similarity to the learned optimal policy. New transition functions are then created by combining elements of the selected pairs. This promotes the exchange of information between transition functions, leading to improved performance. In the mutation stage, individual transition functions undergo small, random changes to introduce diversity and explore new regions of the solution space. The changes are made by randomly adjusting the resulting next states for some state-action pairs. This helps prevent premature convergence and allows for the discovery of novel and effective transition strategies. After completing a predefined number of iterations, the transition function with the highest fitness score is chosen as the optimal representation.
IV Experimental Evaluation
IV-A Experimental Setup
@@
Evaluation metrics in machine unlearning [46] are not applicable to reinforcement unlearning. For instance, in reinforcement unlearning, there are no specific datasets to be forgotten, rendering metrics like “accuracy on forget set” irrelevant. Thus, it is necessary to propose new metrics.
@@ IV-A1 Tasks
The experiments were primarily conducted across three learning tasks: grid world, virtual home and maze explorer. The grid world was developed by us, while the virtual home and maze explorer tasks were sourced from the VirtualHome [40] and MazeExplorer [16], respectively. Although there are well-known RL tasks available, e.g., Gym [15] and Atari [1], they were deemed unsuitable for our research as those RL tasks are designed for single-environment and do not support multiple environments.
@@
Grid World.  The objective for the agent in this task is to navigate towards the predetermined destination. This task mirrors numerous challenges present in real-world autonomous driving and navigation scenarios, with each square akin to a section of road, intersection, or obstacle.
@@
Virtual Home.  It is a multi-agent platform designed to simulate various activities within a household setting. This task mirrors real-world smart home challenges, where homeowners can request the agent to forget certain features of their home.
@@
Maze Explorer.  It is a customizable 3D platform. The objective is to guide an agent through a procedurally generated maze to collect a predetermined number of keys. This task represents scenarios where agents need to learn from visual information, making it relevant to applications such as robotic exploration in intricate environments.
@@
In the three tasks, four available actions include moving up (forward), down (backward), left, and right. Environments are instantiated with predetermined sizes. The instantiation process involves two steps. Firstly, each environment is randomly generated, introducing variability in the placement of obstacles. Subsequently, a manual inspection is carried out to eliminate any instances of “dead locations”, which are inaccessible due to being entirely surrounded by obstacles.
We have also expanded our study to include two additional tasks: recommendation systems and aircraft landing. Recommendation systems are utilized to evaluate privacy risks associated with unlearning, while aircraft landing scenarios serve to assess the safety-critical aspects of our methods.
Recommendation Systems. We employ the MovieLens dataset [18] to simulate a recommendation system using reinforcement learning. In this setup, each user represents a unique training environment. Within this environment, a state is defined by the movies the user has watched and their corresponding ratings. An action corresponds to recommending a specific movie to the user. The reward is determined by the user’s rating of the recommended movie.
Aircraft Landing.  This task simulates an aircraft landing on the ground by avoiding the obstacles with four available actions: moving up, down, left, and right.
@@
While our unlearning methods are evaluated in tasks with discrete state and action spaces, they can also be applied to tasks with continuous state and action spaces. Our approaches are independent of the underlying RL algorithms. To address continuous spaces, one can simply integrate our unlearning techniques with suitable RL algorithms, e.g., [21], designed for such environments.
IV-A2 Comparison Methods
@@
Learning From Scratch (LFS).  This method entails removing the unlearning environment and then retraining the agent from scratch using the remaining environments. However, this method is not a desirable criterion for reinforcement unlearning. We will experimentally show that this approach fails to meet the objectives of reinforcement unlearning as defined in Section III-A.
Non-transferable Learning From Scratch (Non-transfer LFS). To align with the objectives of reinforcement unlearning, we introduce a non-transferable learning-from-scratch approach. This approach is similar to the previously mentioned learning-from-scratch approach. However, a crucial distinction lies in the non-transferable version, which incorporates the non-transferable learning technique [41] to restrict the approach’s generalization ability within the unlearning environments. In this approach, while training an agent, the model owner meticulously stores experience samples acquired from all learning environments, labeling them according to their source environment. When an unlearning request is initiated, the model owner engages in an offline retraining process. Specifically, all the collected experience samples are utilized in retraining the agent. If a sample originates from the unlearning environment, an inverse loss function is applied to minimize the agent’s cumulative reward. Conversely, for samples from other environments, the standard loss function is used to maximize the agent’s overall reward. Denoting the unlearning environment as $\mathcal{M}_{u}=(\mathcal{S}_{u},\mathcal{A}_{u},\mathcal{T}_{u},r)$, the loss functions are defined in Eq. 10.
=$\mathcal{L}=\begin{cases}-\frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max\limits_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)\\ -Q(s_{t},a_{t};\theta))^{2}],&\text{if $s_{t}\in\mathcal{S}_{u}$},\\ \frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max\limits_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)\\ -Q(s_{t},a_{t};\theta))^{2}],&\text{otherwise}.\end{cases}$ (10)
IV-A3 Sample Complexity of Unlearning Methods
@@
There might be a concern about our proposed methods, as they allow the agent to have additional interactions with the unlearning environment. In contrast, both LFS and Non-transfer LFS do not involve further interactions with any environments. However, this additional interaction in our methods does not bring any extra advantages. The purpose of engaging with the unlearning environment is solely to collect experience samples. These experience samples are not required for LFS, and for Non-transfer LFS, they have already been collected. Therefore, the lack of such interactions, does not impact the performance of both LFS and Non-transfer LFS.
IV-A4 Underlying RL Algorithms
The primary RL algorithm employed in this study is the Deep Q-Network (DQN) [27], a well-established value-based method. We also evaluated other RL algorithms, such as Proximal Policy Optimization (PPO) [32] and Deep Deterministic Policy Gradient (DDPG) [22]. However, their performance was found to be inferior to that of DQN (refer to Appendix III). Thus, PPO and DDPG were not included in our experiments.
IV-B Overall Performance
@@
To illustrate the effectiveness of unlearning, we introduce the concept of “forget quality” as a quantitative measure of the strength of forgetting, aligning with the latest machine unlearning evaluation criteria [23] and tailoring it to reinforcement unlearning. We define the “cumulative truth ratio” to gauge this strength, serving as the foundation for the computation of forget quality. Specifically, the cumulative truth ratio $R_{truth}$ can be written as follows.
$R_{truth}=\frac{\sum_{s\in\tau}\frac{1}{|\mathcal{A}_{w}|}\sum_{\hat{a}\in\mathcal{A}_{w}}\pi(\hat{a}|s)}{\sum_{s\in\tau}\pi(\tilde{a}|s)},$ (11)
@@
where $\hat{a}$ represents a wrong action from the wrong action set $\mathcal{A}_{w}$, $\tilde{a}$ is the correct action, and $\tau=((s_{1},a_{1}),\ldots,(s_{N},a_{N}))$ signifies a trajectory in $\mathcal{M}_{u}$. Here, the correct action is defined as the one capable of moving the agent closer to the target compared to other available actions in a given state, while the remaining actions are considered wrong. In particular, the cumulative truth ratio $R_{truth}$ quantifies the ratio between the average probability of selecting wrong actions and the probability of taking the correct action. Thus, a to-be-unlearned agent is expected to achieve a low $R_{truth}$, while an effectively unlearned agent in the unlearning environment should exhibit a high $R_{truth}$, resembling an agent that has never seen $\mathcal{M}_{u}$. For each agent, we obtain $N$ values of $R_{truth}$ each from the first $i$ ($i\in[1,N]$) step(s), and normalize them to serve as an empirical cumulative distribution function (ECDF). A two-sample Kolmogorov-Smirnov (KS) test [24] is conducted between the ECDFs of the unlearned and retained agent to generate a $p$-value to quantify forget quality. A high $p$-value indicates a strong forgetting, implying that the $R_{truth}$ distributions of the unlearned and retained agents are identical.
@@
Figure 5: Forget Quality vs. Model Utility. Larger marker represents more unlearning epochs.
@@
We compute the utility of the unlearned/retained model as the average cumulative rewards of it in the retaining environments. Figure 5 illustrates the trade-off between model utility and forget quality during unlearning in grid world, where a larger marker denotes more unlearning epochs. The decremental RL-based method exhibits consistently high forget quality, while the poisoning-based method displays some fluctuations in forget quality but eventually attains a commendable level. This variance may be attributed to the randomness introduced by the RL algorithm in the poisoning-based approach. During the learning of the poisoning strategy, the method gradually converges, ultimately achieving a favorable result. Furthermore, both methods attain a higher forget quality than before unlearning, indicating the success of the unlearning process.
@@
For model utility, both methods demonstrate robust performance, fluctuating in a narrow range of $(25.075,25.175)$. This performance is comparable to that of the retained models trained by the two baseline methods, i.e., LFS and Non-transfer LFS. A closer examination of model utility reveals a dynamic shift in utility in both methods as the unleanring progresses. This oscillation may arise from two factors. First, the decremental RL-based method aims to erase the agent’s knowledge, potentially leading to over-forgetting [17]. Second, the poisoning-based method involves an intricate interplay between the RL dynamics and the strategic introduction of poison. This interplay can introduce randomness.
(a) Average steps of the four methods
(b) Rewards of the four methods
Figure 6: Four methods in the unlearning environment.
@@
Next, we make detailed comparisons between our methods and the baselines. Figure 6 shows that the unlearning results of the LFS baseline method in all three experimental settings are subpar. The agent’s performance in the unlearning environment remains nearly unchanged before and after unlearning. The reason for this result lies in the agent’s ability to generalize knowledge from other environments and apply it to the unlearning environment, despite never having encountered it before. During training, the agent learns underlying rules and strategies from various environments. For instance, in the grid world setting, the agent acquires knowledge that obstacles should be avoided while collecting the target as quickly as possible. This learned knowledge, even if it was acquired in different environments, enables the agent to still perform well in unseen environments, including the unlearning environment.
@@
(a) The average number of steps before and after unlearning
(b) The average rewards before and after unlearning
Figure 7: The decremental RL-based method in Grid World, where Environment 1 is the unlearning environment.
(a) The average number of steps before and after unlearning
(b) The average rewards before and after unlearning
Figure 8: The poisoning-based method in Grid World, where Environment 1 is the unlearning environment.
However, upon closer comparison between Figures 7 and 8, we can observe slight differences in the performance of the two methods in some remaining environments, such as Environments 18 and 19. In these environments, the poisoning-based method maintains almost unchanged steps and rewards between the pre-unlearning and post-unlearning stages, while the decremental RL-based method does not achieve this. This result suggests that the decremental RL-based method can potentially suffer from the over-unlearning issue to some extent, while the poisoning-based method demonstrates its ability to overcome this issue and retain better performance in the remaining environments after unlearning. These findings highlight the different characteristics and strengths of the two unlearning methods.
(a) The average number of steps before and after unlearning
(b) The average rewards received by the agent before and after unlearning
Figure 9: The decremental RL-based method in Virtual Home, where Environment 1 is the unlearning environment.
(a) The average number of steps before and after unlearning
(b) The average rewards received by the agent before and after unlearning
Figure 10: The decremental RL-based method in Maze Explorer, where Environment 1 is the unlearning environment.
Figures 9 and 10 demonstrate the overall performance of the decremental RL-based method in the context of virtual home and maze explorer, respectively. In the two scenarios, the agent’s behavior exhibits a notable increase in steps taken and a significant decrease in rewards achieved after the unlearning process. The reason for these trends in the two scenarios is rooted in the fundamental nature of reinforcement unlearning. The unlearning process seeks to selectively modify the agent’s behavior to forget specific environments or aspects of its learning history. Thus, the agent must re-explore and adapt to new circumstances, leading to fluctuations in its performance.
(a) The average number of steps before and after unlearning
(b) The average rewards received by the agent before and after unlearning
Figure 11: The poisoning-based method in Virtual Home, where Environment 1 is the unlearning environment.
(a) The average number of steps before and after unlearning
(b) The average rewards received by the agent before and after unlearning
Figure 12: The poisoning-based method in Maze Explorer, where Environment 1 is the unlearning environment.
Figures 11 and 12 provide a comprehensive view of the poisoning-based method’s performance in the virtual home and maze explorer settings, respectively. Remarkably, the performance trend in the two scenarios is similar to that of the decremental RL-based method. This observation reinforces the effectiveness of the poisoning-based approach in reinforcement unlearning, as it consistently achieves the objective of degrading the agent’s performance in the targeted unlearning environment while preserving its capabilities in other environments. The consistent performance trend across different settings shows the method’s versatility and potential applicability in various RL scenarios.
IV-C Hyperparameter Study